# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ IV-A LLMs Tasks
There is also the application of LLMs in ambiguity detection tasks, aimed at detecting ambiguities in natural language requirement documents to improve clarity and reduce misunderstandings. This study primarily aims to address the issue of detecting term ambiguities within the same application domain (where the same term has different meanings in different domains). Although current models generally possess excellent contextual understanding capabilities, this was a common problem in machine learning at that time. This study provides an excellent paradigm for the subsequent application of LLMs in requirements engineering, study demonstrated the transformer-based machine learning models can effectively detect and identify ambiguities in requirement documents, thereby enhancing document clarity and consistency. The framework utilizes BERT and K-means clustering to identify terms used in different contexts within the same application domain or interdisciplinary project requirements documents [57]. In recent two years, more and more researchers use LLMs to help them to evaluate the requirement documentations, quality assessment tasks ensure that the generated requirements and code meet expected quality standards. The application of ChatGPT in user story quality evaluation has shown potential in identifying quality issues, but it requires further optimization and improvement [58]. A similar study use LLM to automatically process the requirement satisfaction assessment, and evaluate whether design elements fully covered by the given requirements, but the the researcher indicated the necessity of further verification and optimization in practical applications [59].
@@ IV-B LLM-based Agents Tasks
Currently the application of LLM-based agents in the requirement engineering is till quite nascent, but there are some useful researches to help us to see the potential possibility. LLM-based agents bring both efficiency and accuracy for tasks like requirement elicitation, classification, generation, and verification. Compared to traditional LLMs, these systems exhibit higher levels of automation and precision through task division and collaboration. The application of multi-agent systems in semi-structured document generation has shown significant effectiveness. In [60], a multi-agent framework is introduced that combines semantic recognition, information retrieval, and content generation tasks to streamline the creation and management of semi-structured documents in the public administration domain. The proposed framework involves three main types of agents: Semantics Identification Agent, Information Retrieval Agent, and Content Generation Agent. By avoiding the overhead of a single model, each agent is assigned a specific task with minimal user intervention, following the designed framework and workflow.
@@
Furthermore, in generating safety requirements for autonomous driving, LLM-based agents have shown unique advantages by introducing multimodal capabilities. The system employs LLMs as automated agents to generate and refine safety requirements with minimal human intervention until the verification stage, which is unattainable with only LLMs. [62] describes an LLM prototype integrated into the existing Hazard Analysis and Risk Assessment (HARA) process, significantly enhancing efficiency by automatically generating specific safety-related requirements. The study through three design iterations progressively improved the LLM prototype’s efficiency by completing within a day compared to months manually. In agile software development, the quality of user stories directly impacts the development cycle and the realization of customer expectations. [63] demonstrates the successful application of the ALAS system in six agile teams at the Austrian Post Group IT. The ALAS system significantly improved the clarity, comprehensibility, and alignment with business objectives of user stories through automated analysis and enhancement. The entire agent framework allows the model to perform specific roles in the Agile development process, the study results indicated that the ALAS-improved user stories received high satisfaction ratings from team members.
@@ IV-C Analysis
The application of LLM-based agents in requirement engineering has demonstrated significant efficiency improvements and quality assurance. Through multi-agent collaboration and automated processing, these systems not only reduce manual intervention but also enhance the accuracy and consistency of requirement generation and verification. We can see that the tasks of LLM-based agents are no longer limited to simply generating requirements or filling in the gaps in descriptions. Instead, they involve the implementation of an automated process, with the generation of requirement documents being just one part of it, integrating LLM into agents enhances the overall system’s natural language processing and reasoning capabilities. In the real-world application, many tasks can no longer be accomplished by simple LLMs alone, especially for high-level software design. The emergence of LLM-based agents addresses this issue through a multi-agent collaborative system centered around LLMs, these agents continuously analyze and refine the deficiencies in the requirement documents, this is might be the main application trend of LLM-based agents in requirements engineering in the future.
Figure 4: Illustration of Comparison Framework Between LLM-based Agent and LLM in User Story Refinement
@@ IV-D Benchmarks
Requirement engineering, unlike tasks such as bug fixing and code generation, does not have an abundance of public datasets available, such as HumanEval which commonly used for code generation assessment. Most training datasets for models in requirement engineering are self-collected by the authors and not all of them are open-sourced on Huggingface, resulting in a limited amount of dataset in requirement engineering. For instance, some papers do not mention a specific benchmark dataset but instead focus on practical examples and case studies to demonstrate the effectiveness of proposed prompt patterns [55]. The researcher Let actual developers and requirements engineers use the generated requirements documents and code to evaluate its accuracy, usability, and completeness. User feedback will be collected to further improve and optimize the prompt mode.
@@
TABLE V: Evaluation Metrics in Requirement Engineering and Documentation
@@
[…]
 [49]] PROMISE NFR Precision (P), Recall (R), F1-score (F1),
@@ IV-E Evaluation Metrics
In the field of requirement engineering, LLMs and LLM-based agents are evaluated using various metrics. These metrics not only include traditional indicators such as precision, recall, and F1 score but also more specific indicators tailored to the unique nature of requirement engineering. Through these evaluations, we can see how these models are assessed and how they are changing the practice of requirement engineering. The specific evaluation metrics are detailed in Table V. In [51], while precision and recall are fundamental for evaluating the effectiveness of information retrieval, additional evaluations of clarity, consistency, and compliance are included, which are crucial quality indicators in requirement engineering. This multidimensional evaluation method not only measures the operational performance of LLMs but also examines their ability to maintain the quality of requirement specifications. Through this approach, LLMs have demonstrated their value in automating and optimizing the requirement elicitation process, enhancing both efficiency and the reliability of results. The paper  [52] use the Likert Scale to measure the quality of generated specifications, the specification will be scored by its Unambiguous, Understandable, Conciseness, etc. The Likert Scale will be scored from 1 to 5 of agreements.
For agent-based LLMs, as demonstrated in [63], the evaluation extends to assessing the independence and negotiability of the agents, elevating their functionality to a new level. These agents provide technical solutions and also interact with users, autonomously adjusting to meet specific project needs, thus resembling collaborative partners. This capability makes LLM-based agents valuable in requirement engineering in the requirement management and decision optimization, also highlight that LLMs typically focus on improving the accuracy and efficiency of specific tasks, while LLM-based agents exhibit higher capabilities in autonomy and adaptability.
In Table V, we can see that the application of LLMs in requirements engineering typically requires common metrics such as F1 Score to evaluate model’s performance. However, for LLM-based agents, the evaluation focus shifts from the performance of requirement document generation to the quality of the final product. Therefore, evaluation metrics emphasize user satisfaction, such as pass rate, feedback, etc. Essentially, LLM-based agents still leverage LLMs themselves to achieve higher-level development, and depends pretty much on the nature of the task. In summary we can conclude that, the characteristics of agent models both reflect their complex decision-making and learning abilities also reveal their potential advantages in collaboration with human or other tools to provide higher scalability and flexibility design. This phenomenon implies the potential opportunities that the methods of requirement elicitation and processing in future software development will become more efficient, precise, and continuous refine to better align with stakeholder’s needs by using LLM-based agents.
@@ V-A LLMs Tasks
Large language models have optimized various tasks in code generation and software development through automation and reasoning, covering areas such as code generation, debugging, code comprehension, code completion, code transformation, and multi-turn interactive code generation. The primary method is generating executable code from natural language descriptions, where models utilize previously learnt code snippets or apply few-shot learning to better understand user requirements. Nowadays the AI tools integrates deeply with IDEs like Visual Studio Code [^2] https://code.visualstudio.com/ and JetBrains [^3] https://www.jetbrains.com/ to enhance code writing and translation tasks such as OpenAI’s Codex model [67]. Codex fine-tuned on public code from GitHub, demonstrate the capability to generate Python functions from doc-strings also outperformed other similar models on the HumanEval benchmark.
@@
Additionally, the application of LLMs in improving programming efficiency has garnered widespread attention, the tools like GitHub Copilot which integrating OpenAI’s Codex model, provide real-time code completion and suggestions during coding. According to [72], researchers present a controlled experiment with the Github Copilot, the result demonstrated that with developers completing HTTP server tasks 55.8% faster when using Copilot. Another similar study also using LLM to be the programmer tools, in [73], researchers introduced the INCODER model which capable of both program synthesis and editing. By leveraging bidirectional context, the model performs well in both single-line and multi-line code filling tasks, providing developers with smarter code editing tools. This real-time code generation and completion functionality not only improves programming efficiency but also reduce the burden on developers, allowing them to focus on higher-level design which is a common problem in software development where substantial workforce and time are wasted on tedious coding tasks.
@@
The multi-turn program synthesis tasks represent a significant breakthrough for LLMs in handling complex programming tasks, in  [74], researchers introduced the CODEGEN model, which iteratively generates programs through multiple interactions, significantly improving program synthesis quality and making the development process more efficient and accurate. By gradually generating and continuously optimizing code at each interaction, LLMs can better understand user intent and generate more precise and optimized code. In the experiments, comparisons were made with the Codex model, which was considered state-of-the-art in code generation at the time. CODEGEN-MONO 2.7B outperformed the Codex model of equivalent outcome in pass@k metrics for both k=1 and k=10. Furthermore, CODEGEN-MONO 16.1B exhibited performance that was comparable to or better than the best Codex model on certain metrics, further demonstrating its SOTA performance in the code generation. By iteratively generating and optimizing code, LLMs continuously improve their output quality. In [75], researchers proposed the Cycle framework, which enhances the self-improvement capability of code language models by learning from execution feedback, improving code generation performance by 63.5% on multiple benchmark datasets. Although Cycle has a certain degree of autonomy, its decision-making and planning capabilities are mainly limited to code generation and improvement tasks without overall planning, and the execution sequence is completely followed a fixed pattern, so it’s better to classified as an advanced LLM application.

### v2

@@ IV-A LLMs Tasks
Ambiguity Detection in Requirement Documents. There is also the application of LLMs in ambiguity detection tasks, aimed at detecting ambiguities in natural language requirement documents to improve clarity and reduce misunderstandings. This study primarily aims to address the issue of detecting term ambiguities within the same application domain (where the same term has different meanings in different domains). Although current models generally possess excellent contextual understanding capabilities, this was a common problem in machine learning at that time. This study provides an excellent paradigm for the subsequent application of LLMs in requirements engineering, study demonstrated the transformer-based machine learning models can effectively detect and identify ambiguities in requirement documents, thereby enhancing document clarity and consistency. The framework utilizes BERT and K-means clustering to identify terms used in different contexts within the same application domain or interdisciplinary project requirements documents [72].
Requirement Quality Assessment and Structured Evaluation. In recent two years, more and more researchers use LLMs to help them to evaluate the requirement documentations, quality assessment tasks ensure that the generated requirements and code meet expected quality standards. The application of ChatGPT in user story quality evaluation has shown potential in identifying quality issues, but it requires further optimization and improvement [73]. A similar study use LLM to automatically process the requirement satisfaction assessment, and evaluate whether design elements fully covered by the given requirements, but the the researcher indicated the necessity of further verification and optimization in practical applications [74]. Further extending this direction, Ronanki et al. [75] systematically compared ChatGPT-generated requirements with those written by RE professionals across seven quality attributes. ChatGPT excelled in understandability and correctness, but scored lower in feasibility and unambiguity, especially when domain-specific nuances were required. Complementing this, Yeow et al. [76] examined GPT-3.5’s ability to generate elicitation questions across multiple domains—including banking, healthcare, education, and retail. Through a combination of readability metrics and expert scoring, they found the questions to be generally clear, relevant, and complete. The model maintained consistent quality across domains, demonstrating potential to assist early-stage RE, though issues like specificity and depth still require refinement. Building on this progression, Wei [77] proposed a ”progressive prompting” strategy that aligns with the waterfall model. Their system guided GPT-3.5 through structured stages—from requirements to use cases, designs, and code—using iterative prompts and knowledge bases. Applied to a case study, the approach improved traceability and explainability, positioning LLMs as semi-agentic partners in structured RE and software design workflows.
@@ IV-B LLM-based Agents Tasks
Multi-Agent Collaboration and Autonomous Iteration. Currently the application of LLM-based agents in the requirement engineering is till quite nascent, but there are some useful researches to help us to see the potential possibility. LLM-based agents bring both efficiency and accuracy for tasks like requirement elicitation, classification, generation, and verification. Compared to traditional LLMs, these systems exhibit higher levels of automation and precision through task division and collaboration. The application of multi-agent systems in semi-structured document generation has shown significant effectiveness. In [78], a multi-agent framework is introduced that combines semantic recognition, information retrieval, and content generation tasks to streamline the creation and management of semi-structured documents in the public administration domain. The proposed framework involves three main types of agents: Semantics Identification Agent, Information Retrieval Agent, and Content Generation Agent. By avoiding the overhead of a single model, each agent is assigned a specific task with minimal user intervention, following the designed framework and workflow.
@@
Safety Requirements and Agile User Stories. Furthermore, in generating safety requirements for autonomous driving, LLM-based agents have shown unique advantages by introducing multimodal capabilities. The system employs LLMs as automated agents to generate and refine safety requirements with minimal human intervention until the verification stage, which is unattainable with only LLMs. [80] describes an LLM prototype integrated into the existing Hazard Analysis and Risk Assessment (HARA) process, significantly enhancing efficiency by automatically generating specific safety-related requirements. The study through three design iterations progressively improved the LLM prototype’s efficiency by completing within a day compared to months manually. In agile software development, the quality of user stories directly impacts the development cycle and the realization of customer expectations. [81] demonstrates the successful application of the ALAS system in six agile teams at the Austrian Post Group IT.customized bjectives of user stories through automated analysis and enhancement. The entire agent framework allows the model to perform specific roles in the Agile development process, the study results indicated that the ALAS-improved user stories received high satisfaction ratings from team members.
Requirement Compliance and Workflow Automation. Regulatory compliance challenges are addressed through advanced agent architectures. Masoudifard et al. [82] integrate Graph-RAG [83] with reasoning techniques (Chain-of-Thought, Tree-of-Thought) to validate requirements against regulations. Tested in financial and aerospace systems (e.g., NASA’s X-38), their framework achieves 87.93% F1-score in accuracy, demonstrating scalability for mission-critical contexts. A comprehensive approach is exemplified by Sami et al. [84], who simulate end-to-end requirement workflows using role-specific agents (Product Owner, QA, Developer, Manager). This multi-agent system automates user story generation, quality assessment, feasibility review, and prioritization, achieving coherence and efficiency comparable to human teams. Evaluations highlight GPT-3.5’s balance of performance and Mixtral-8B’s speed, while practitioners emphasize its potential to enhance human-AI collaboration. These advancements underscore LLM-based agents’ versatility across requirement engineering—from granular tasks like compliance checking to holistic pipeline automation—while emphasizing the enduring role of human expertise in validation and decision-making.
@@ IV-C Analysis
The application of LLM-based agents in requirement engineering has demonstrated significant efficiency improvements and quality assurance [67, 80]. Through multi-agent collaboration and automated processing, these systems not only reduce manual intervention but also enhance the accuracy and consistency of requirement generation and verification [78]. We can see that the tasks of LLM-based agents are no longer limited to simply generating requirements or filling in the gaps in descriptions. Instead, they involve the implementation of an automated process, with the generation of requirement documents being just one part of it, integrating LLM into agents enhances the overall system’s natural language processing and reasoning capabilities [85]. In the real-world application, many tasks can no longer be accomplished by simple LLMs alone, especially for high-level software design. The emergence of LLM-based agents addresses this issue through a multi-agent collaborative system centered around LLMs, these agents continuously analyze and refine the deficiencies in the requirement documents, this is might be the main application trend of LLM-based agents in requirements engineering in the future.
Figure 5: Illustration of Comparison Framework Between LLM-based Agent and LLM in User Story Refinement.
@@ IV-D Benchmarks
Requirement engineering, unlike tasks such as bug fixing and code generation, does not have an abundance of public datasets available, such as HumanEval which commonly used for code generation assessment. Most training datasets for models in requirement engineering are self-collected by the authors and not all of them are open-sourced on Huggingface, resulting in a limited amount of dataset in requirement engineering. For instance, some papers do not mention a specific benchmark dataset but instead focus on practical examples and case studies to demonstrate the effectiveness of proposed prompt patterns [70]. Some researchers involve actual developers and requirements engineers to evaluate the accuracy, usability, and completeness of generated requirements and code artifacts through practical usage and feedback.
@@
A closer examination reveals a clear shift in benchmark design accompanying the transition from LLM-based RE methods to agent-based systems. Benchmarks in LLM studies tend to focus on isolated tasks, such as requirement classification, term extraction, or user story generation. These datasets are often small-scale, domain-specific, and emphasize text-level correctness or completeness—frequently evaluated through human ratings or token-level metrics. The scope is typically narrow, targeting single-turn input-output behavior of the model. In contrast, LLM-based agent research introduces a new benchmark paradigm: datasets are increasingly designed to simulate realistic RE workflows, involving multi-turn reasoning, multi-agent collaboration, and cross-document traceability. For instance, recent benchmarks such as SRS Broker and SRS Aero [82] are built around regulatory compliance, containing rich interdependencies among requirements and standards. Evaluation expands beyond accuracy to include reasoning traceability, compliance violation detection, and structured knowledge retrieval. Moreover, agent-based benchmarks often incorporate role-driven interaction settings, as seen in Sami et al. [84], where Product Owners, Developers, QA Engineers, and Project Managers coordinate through LLM agents. Evaluation thus moves toward semantic alignment, system response time, prioritization quality, and multi-perspective expert judgment, often grounded in international standards such as ISO/IEC 29148 or the INVEST framework. This evolution marks a significant step from task-oriented benchmarking to scenario- and collaboration-driven evaluation. This progression indicates that as LLMs evolve from passive tools to autonomous agents, benchmark datasets must likewise evolve—from measuring isolated task performance to evaluating contextual understanding, dynamic decision-making, and agentic collaboration in real-world RE environments.
TABLE VI: Evaluation Metrics in Requirement Engineering and Documentation
@@
PTC-B $F\beta$, Mean Average Precision (MAP) No
 [61] PROMISE NFR Precision (P), Recall (R), F1-score (F1),
@@
 [69] 15 UML models, 168 English specs Validity@K, Correctness@K, McNemar’s Test No
 [65] 7 RE documents RUST framework, Likert scale No
 [63] PROMISE, Dronology, ReqView, Leeds Library, WASP $F\beta$, Accuracy, Recall No
 [75] 36 requirements (6 AI, 30 human) 7 RE Quality Attributes No
 [76] 4 domains survey and QA Flesch Reading Ease, FK Grade Level, Clarity,
Relevance, Completeness No
 [77] SuperFrog Scheduler case study Accuracy, Adherence to Requirements,
Practical Usability No
@@ IV-E Evaluation Metrics
The evaluation of LLMs and LLM-based agents in requirement engineering spans from classical NLP performance metrics to RE-specific quality attributes and agent-oriented behavioral indicators. As summarized in Table VI, evaluation methods are closely tied to the task nature—whether static classification, document generation, or interactive system behavior.
For traditional LLM applications, especially in classification and extraction tasks (e.g., functional vs. non-functional requirements), common metrics include precision, recall, and F1 score, with several studies applying class-weighted Fβ scores to emphasize recall in imbalanced datasets [63, 62]. In generative scenarios, such as user story or specification generation, human-centric evaluation becomes prominent. For instance, the RUST framework [65] scores GPT-4’s outputs on Readability, Understandability, Specifiability, and Technicality using Likert scales. Similarly, SRS evaluations in [66] employ 1–5 Likert ratings for Conciseness, Consistency, and Completeness, while [75] compares ChatGPT-generated requirements with expert-authored ones using seven RE-specific quality attributes including Correctness, Unambiguity, and Feasibility. Domain-specific ambiguity detection models evaluate semantic divergence using contextual clustering [72], and completeness assessments for masked terms use precision-oriented metrics [71].
As models shift from passive language tools to autonomous agents, the scope of evaluation expands significantly. In collaborative and role-specific agent settings [84], evaluation considers semantic similarity (via sentence embeddings), response time, and alignment with standards like ISO/IEC 29148 and the INVEST criteria for user stories. The assessment includes human feedback across four real-world roles: Product Owner, Developer, QA, and Project Manager, offering multi-perspective validation. In compliance-oriented tasks such as those tackled by Graph-RAG agents [82], evaluation focuses on compliance classification accuracy, F1 scores for non-compliance detection, and reasoning interpretability under various prompting strategies (e.g., Chain-of-Thought, Tree-of-Thought). These methods evaluate not only output correctness but also the explainability and robustness of agent decision paths.
Agent-driven frameworks like AISD [79] adopt system-level metrics such as use case pass rate, code consistency, and human-LLM iteration count. These reflect the agents’ ability to refine and adapt outputs over multiple feedback loops. Similarly, in safety-critical environments [80], evaluation involves task duration reduction and expert validation of safety requirement adequacy across design iterations. This transition from LLM to agent evaluation represents a fundamental shift: from static, single-turn outputs measured by token-level correctness to dynamic, multi-role collaboration assessed by behavioral robustness, human satisfaction, and contextual traceability. While LLMs excel in task precision, agent-based systems are evaluated for their autonomy, adaptability, and collaborative fluency. This shift implies that future RE systems must embrace richer evaluation schemas, integrating quantitative performance with qualitative feedback and scenario-driven benchmarks, to capture the full spectrum of agentic capability in evolving software engineering workflows.
@@ V-A LLMs Tasks
Code Generation and Translation from Natural Language. Large language models have optimized various tasks in code generation and software development through automation and reasoning, covering areas such as code generation, debugging, code comprehension, code completion, code transformation, and multi-turn interactive code generation. The primary method is generating executable code from natural language descriptions, where models utilize previously learnt code snippets or apply few-shot learning to better understand user requirements. Nowadays the AI tools integrates deeply with IDEs like Visual Studio Code [^2] https://code.visualstudio.com/ and JetBrains [^3] https://www.jetbrains.com/ to enhance code writing and translation tasks such as OpenAI’s Codex model [88]. Codex fine-tuned on public code from GitHub, demonstrate the capability to generate Python functions from doc-strings also outperformed other similar models on the HumanEval benchmark.
@@
Real-Time Code Completion and Editing Tools. Additionally, the application of LLMs in improving programming efficiency has garnered widespread attention, the tools like GitHub Copilot which integrating OpenAI’s Codex model, provide real-time code completion and suggestions during coding. According to [93], researchers present a controlled experiment with the Github Copilot, the result demonstrated that with developers completing HTTP server tasks 55.8% faster when using Copilot. Another similar study also using LLM to be the programmer tools, in [94], researchers introduced the INCODER model which capable of both program synthesis and editing. By leveraging bidirectional context, the model performs well in both single-line and multi-line code filling tasks, providing developers with smarter code editing tools. This real-time code generation and completion functionality not only improves programming efficiency but also reduce the burden on developers, allowing them to focus on higher-level design which is a common problem in software development where substantial workforce and time are wasted on tedious coding tasks.
@@
Multi-Turn Program Synthesis with Feedback. The multi-turn program synthesis tasks represent a significant breakthrough for LLMs in handling complex programming tasks, in  [95], researchers introduced the CODEGEN model, which iteratively generates programs through multiple interactions, significantly improving program synthesis quality and making the development process more efficient and accurate. By gradually generating and continuously optimizing code at each interaction, LLMs can better understand user intent and generate more precise and optimized code. In the experiments, comparisons were made with the Codex model, which was considered state-of-the-art in code generation at the time. CODEGEN-MONO 2.7B outperformed the Codex model of equivalent outcome in pass@k metrics for both k=1 and k=10. Furthermore, CODEGEN-MONO 16.1B exhibited performance that was comparable to or better than the best Codex model on certain metrics, further demonstrating its SOTA performance in the code generation. By iteratively generating and optimizing code, LLMs continuously improve their output quality. In [96], researchers proposed the Cycle framework, which enhances the self-improvement capability of code language models by learning from execution feedback, improving code generation performance by 63.5% on multiple benchmark datasets. Although Cycle has a certain degree of autonomy, its decision-making and planning capabilities are mainly limited to code generation and improvement tasks without overall planning, and the execution sequence is completely followed a fixed pattern, so it’s better to classified as an advanced LLM application. Another representative multi-turn code generation framework is TICODER [97], which integrates test-driven interaction to disambiguate user intent and improve LLM code generation quality. Rather than directly surfacing candidate code completions, TICODER generates tests based on the user’s natural language intent and asks users to validate them. Based on user responses (PASS, FAIL, or OUTPUT), it incrementally prunes invalid code suggestions and surfaces a refined, ranked list of consistent completions. A user study with 15 programmers showed that TICODER significantly improved correctness (84% vs. 40% in baseline), reduced cognitive load (mean TLX score 28 vs. 45), and did not increase task time. Furthermore, in large-scale benchmark evaluations on MBPP and HumanEval, TICODER showed average absolute improvements of 45.73% in pass@1 accuracy within five user interactions across diverse LLMs (including GPT-3.5, GPT-4, CodeGen, and Codex). These results confirm LLM framework’s strength as a semi-structured, test-guided workflow that supports LLMs in interactive, high-precision code generation. Another framework that embraces multi-phase refinement is AlphaCodium [98], which proposes a test-driven, flow-based approach specifically optimized for solving complex competitive programming tasks. Unlike one-shot prompting or static few-shot examples, AlphaCodium introduces a structured workflow that includes (1) pre-processing through problem reflection and reasoning over public tests, and (2) iterative refinement via both public and AI-generated test cases. The process incorporates design patterns such as semantic bullet-point analysis, modular code generation, and structured YAML outputs. A key insight from the study is that generating diverse test cases is often easier and more reliable for LLMs than producing complete correct code solutions. AlphaCodium thus leverages this by iteratively running candidate programs against a curated test suite and refining the outputs until convergence. Another iterative refinement method is RepoCoder [99], which addresses repository-level code completion through multi-round retrieval-augmented generation. RepoCoder retrieves relevant code snippets from the same repository based on unfinished code, generates candidate completions using an LLM (e.g., GPT-3.5), and reuses generated code to iteratively refine retrieval. Evaluated on the RepoEval benchmark with function-, API-, and line-level tasks, it achieved up to 10% improvements in exact match and functional correctness over one-shot and in-file methods. The study demonstrates that iterative retrieval and synthesis significantly improve contextual grounding and code quality without requiring agentic coordination.
LLM-Orchestrated Algorithm Synthesis with Verifiers. While traditional LLM-based code generation excels at translating natural language into executable programs, it struggles with algorithm synthesis tasks that require not just implementation but also the inference of the correct algorithmic idea. To address this, ALGO [100] introduces a novel framework where LLMs are used not only as code generators but also as oracle verifiers. ALGO consists of two modules: a verifier that generates a brute-force but correct oracle using LLMs, and a coder that synthesizes efficient solutions guided by the verifier’s outputs. The oracle is created by prompting LLMs to exhaustively enumerate the search space and is used to validate candidate programs by comparing outputs on a batch of generated inputs. Experimental results show that ALGO significantly boosts synthesis performance across various LLM coders: achieving an 8× increase in one-shot pass rate over Codex and a 2.6× gain over CodeT.
While LLMs have demonstrated impressive performance across various static and feedback-driven code generation tasks, their ability to autonomously plan, reason across stages, or coordinate within larger development workflows remains limited. This gap has motivated recent research into LLM-based agents that embody higher-level autonomy and collaboration capabilities, as discussed in the following section.